OpenStudy (kkutie7):

I feel like this second part of this problem might be able to be solved by math (conversions and such) so please help me if you can.

OpenStudy (kkutie7):

From the Faradaic current flowing in the cell, calculate the moles of mercury that plated onto the graphite electrode tip. From this, calculate the approximate thickness of the mercury film on the tip of the electrode.

OpenStudy (kkutie7):

deposition time=1min. average cell current during deposition= 1mA density of mercury= 13.54 g/mL diameter of electrode=4.0 mm Faradaic constant= 96485 c/ mole of electron

OpenStudy (kkutie7):

\[1mA*\frac{1A}{1000mA}*60s=0.06C*\frac{1 mole of electron}{96485c}*\] \[\frac{1moleHg}{2moleof electron}=3.1092915*10^{-7}moleHg\]

OpenStudy (kkutie7):

this is as far as i can go

OpenStudy (jiteshmeghwal9):

\[\frac{I×t}{96485}= \frac{W}{E}\]

OpenStudy (jiteshmeghwal9):

W= weight of Mercury deposited E=equivalent weight

OpenStudy (kkutie7):

mind working it a little more out i'm slightly confused

OpenStudy (agent0smith):

Looks like you have moles, and your calculation looks reasonable to me. With moles, you can find mass, and then from mass (and density), you can find the volume of the Hg. You'd need some other dimension to find the thickness, though.

OpenStudy (jiteshmeghwal9):

\[\frac{10^{-3} \times 60}{96485}=n \times 2\]where n is the number of moles of mercury & 2 is the n-factor u may write \[\frac{W}{E}=n \times \text{n-factor}\]

OpenStudy (kkutie7):

are you solving for the moles of mercury? I don't mean to be a pain but I need help with the thickness and I don't see that your equation works for that. It does look close to what I did to find moles though

OpenStudy (agent0smith):

@Kkutie7 see what I said above. Convert moles of Hg to mass, and then use density = mass/volume.

OpenStudy (kkutie7):

oh I must have missed that

OpenStudy (agent0smith):

Oh and you do have the diameter.

OpenStudy (kkutie7):

yupp of the electrode

OpenStudy (kkutie7):

thickness is volume/area?

OpenStudy (agent0smith):

Yeah that should work.